OpenStudy (anonymous):

If two dies are tossed together,what is the probability of having equal outcomes

OpenStudy (anonymous):

what do you mean by equal outcomes.

OpenStudy (anonymous):

having the same outcomes like 11 22 33 44 55 66

OpenStudy (anonymous):

ok, so you want to find P(1 AND 1) +P(2 AND 2) + .... P(1 AND 1) = P(1) * P(1) = (1/6)*(1/6) so in the end i think you have 6[(1/6)*(1/6)]